OpenStudy (anonymous):

Divergence of a constant vector. I tried calculating the divergence of a constant vector but I get it equal to zero. The sample solution in my book says the following:

OpenStudy (perl):

what does your book says?

OpenStudy (anonymous):

I'm typing it now, it will take a bit to type the formulas.

OpenStudy (anonymous):

\[( \nabla \cdot B_0 ) r = B_0 \cdot \nabla r\]

OpenStudy (anonymous):

Where B0 is a constant vector.

OpenStudy (jtvatsim):

what is r?

OpenStudy (anonymous):

Another vector, I suppose the radius vector.

OpenStudy (anonymous):

The divergence of a constant vector will be 0.

OpenStudy (anonymous):

So then is the textbook wrong?

OpenStudy (anonymous):

This would make more sense:\[( \nabla \cdot B_0 r) = B_0 \cdot \nabla r\]

OpenStudy (anonymous):

Yes, indeed.

OpenStudy (anonymous):

Hmmm, maybe it's constant with respect to something else?

OpenStudy (anonymous):

I think the book is mistaken, or we are talking about something far more complicated

OpenStudy (anonymous):

Let me copy the whole problem. Maybe I'm missing something else.

OpenStudy (anonymous):

The connection between the magnetic field B and vector potential A is \[ B = \nabla \times A\] What is the magnetic field if \[A = \frac{ 1 }{ 2 }B_0 \times r\] where B0 is a constant vector?

OpenStudy (anonymous):

So, the solution says:\[B = \nabla \times ( \frac{ 1 }{ 2 }B_0 \times r) = \frac{ 1 }{ 2 }B_0 ( \nabla \cdot r) - \frac{ 1 }{ 2 } ( \nabla \cdot B_0 ) r \]

OpenStudy (anonymous):

And then they claim that \[( \nabla \cdot B_0 ) r = B_0 \cdot \nabla r\]

OpenStudy (anonymous):

Does it make any sense now, @wio ?

OpenStudy (anonymous):

Maybe \(B_0\) isn't constant then? To be honest I find it a bit strange.

OpenStudy (anonymous):

Like, maybe it's constant for time, but not for position

OpenStudy (anonymous):

So \(B_0 = f(x,y,z)\) but \(d/dt B_0 = 0\)

OpenStudy (anonymous):

Right, I see. Well, I guess I should ask the teacher, then. Thank you for taking the time :)

OpenStudy (anonymous):

Okay, one way to think about this is: \[ (B_0)_xr_1+(B_0)_yr_2 = (B_0)_1r_x + (B_0)_2r_y \]

OpenStudy (anonymous):

I'm not sure I understand what you did. Why are the components of B0 and r different - once expressed through x and y, and in the other case - through 1 and 2?